OpenStudy (anonymous):

Find the equation of a line with the two points (-10,6) and (-8,7) Is it A)y+6=1/2(x+10) B)y+10=1/2(x-6) C)y-6=1/2(x+10) or D)y-6=1/2(x-10)

OpenStudy (anonymous):

Use the formula \[ \frac { y-y_1}{x-x_1}= \frac { y_2-y_1}{x_2-x_1}=m\\ y-y_1 =m(x-x_1) \]

OpenStudy (anonymous):

\[ m=\frac { 7-6}{-8+10}=\frac 1 2 \]

OpenStudy (anonymous):

(-10,6) and (-8,7) \[ y-6= \frac 1 2(x+10) \\ \]

OpenStudy (anonymous):

Thank you :)